OpenStudy (anonymous):

find the derivative of y=ln2* log2 x

OpenStudy (anonymous):

The books said the answer is 1/x but I got 1/2xln2 as my answer?

zepdrix (zepdrix):

Change of base formula:\[\large\rm \log_2x=\frac{\ln x}{\ln 2}\]

zepdrix (zepdrix):

Do you see how the ln(2)'s will cancel out?

OpenStudy (anonymous):

wouldn't it be x/2 then?

OpenStudy (anonymous):

oh wait a minute... ln 2* 1/x ln2= 1/x right?

zepdrix (zepdrix):

I would recommend canceling them out `before` differentiating :) Maybe will be easier to understand that way.\[\large\rm y=\ln2\cdot\color{orangered}{\log_2x}\]\[\large\rm y=\ln2\cdot\color{orangered}{\frac{\ln x}{\ln2}}\]\[\large\rm y=\ln x\]But yes, your way works fine as well :)

OpenStudy (anonymous):

Thanks for explaining :)